OpenStudy (anonymous):

Integrate (sinx)^4(cosx)^2

OpenStudy (anonymous):

wolfram aplha i got pretty far n not enough space to see where i did the mistake.

OpenStudy (anonymous):

its REALLY UGLY you set into cos^3(2x) and have to combine terms i got to this point 3/8+(-7/8cos(2x))+(5/8cos(2x))+(-1/8cos^3(2x))

OpenStudy (anonymous):

3/8+(-7/8cos(2x))+(5/8cos^2(2x))+(-1/8cos^3(2x))

OpenStudy (anonymous):

-1/8cos^3(2x)=-1/8(cos(2x)(1-sin^2(2x)) -1/8cos(2x)+1/8cos(2x)sin^2(2x)

OpenStudy (anonymous):

cos^2(2x)=(1/2+1/2cos(4x))

OpenStudy (anonymous):

3/8+(-7/8cos(2x))+(5/8(1/2+1/2cos(4x)))+(-1/8cos(2x)+1/8cos(2x)sin^2(2x))

OpenStudy (anonymous):

3/8+5/16=11/16 -cos(2x) +1/2cos(4x) 1/8cos(2x)sin^2(2x)\[1/8\int\limits_{}^{}\cos^2(2x)\sin^2(2x)dx\]+\[\int\limits_{}^{}1/2\cos(4x)dx\]+\[-\int\limits_{}^{}\cos(2x)dx\]+\[\int\limits_{}^{}11/16dx\] \[\int\limits_{}^{}1/8\cos(2x)\sin^2(2x)=1/48(\sin^3(2x))\] \[\int\limits_{}^{}1/2\cos(4x)=1/8\sin(4x)\] \[\int\limits_{}^{}-\cos(2x)=-1/2\sin(2x)\] \[\int\limits_{}^{}11/16=11/16+C\] 1/48sin^3(2x)+1/8sin(4x)-1/2sin(2x)+11/16x+C

OpenStudy (anonymous):

which is not whats on wolfram ha anyways